OpenStudy (anonymous):

An electrician cuts a 26 ft piece of wire into two pieces. One piece is 2 ft longer than the other. How long are the pieces?

OpenStudy (anonymous):

Let x,y are the two pieces. x+y=26

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

and let x is greater than 2 ft than y

OpenStudy (anonymous):

greater means +

OpenStudy (anonymous):

x=y+2

OpenStudy (anonymous):

ok than what

OpenStudy (anonymous):

put the value of in 1st equation. 2y+2=26 2y=24 y=12 x=y+2 x=14

OpenStudy (anonymous):

so shorter piece 12 & longer piece 14

OpenStudy (anonymous):

yeah